The following are the daily returns for both the overall market and for Omega Inc. What is the cumulative abnormal return on Omega, Inc., stock for these 5 days?

Date R(m) Omega
24-Nov 1.20 % 1.20 %
25-Nov 0.40 % 0.10 %
26-Nov 0.10 % 1.00 %
27-Nov 0.40 % 0.50 %
28-Nov 0.50 % 0.70 %

Group of answer choices

–0.1 percent

0.2 percent

–0.5 percent

–0.8 percent

0.5 percent
Answer #1

Ans:- In this question, we need to find the Cumulative abnormal return on the Omega Inc.stock for 5 days.

Cumulative abnormal return is nothing but the financial term which is used in the market to describe investment value. It basically describes the relationship between the expected return on the stock with the overall market performances and the actual value of the stock.

Cumulative abnormal return is calculated by subtracting the market return from the individual return on a stock.

Date Return on Market Return on Omega Inc. Abnormal Return
24 Nov 1.20% 1.20% 1.20 - 1.20 = 0%
25 Nov -0.40% 0.10% 0.10 - (-40) = 0.5%
26 Nov -0.10% 1.0% 1.0 - (-0.10) =1.1%
27 Nov 0.40% -0.50% -0.5 - 0.40 = -0.9%
28 Nov 0.50% -0.70% -0.7 - 0.50 = -1.2%
Total = -0.5%

Cumulative abnormal return is 0 + 0.5 + 1.1 -0.9 -1.2 = -0.5% is the required answer.

Note:- Please give thumbs up if this answer helps you.